# Real $K$-Theory for $C^{*}$-Algebras: just the facts October 22, 2025

### v2

@@ 8. Bott Periodicity
We now recap the principal properties of $KO_{*}$ for future use. Let $\mathscr{C}^{*}_{real}$ denote the category of real $C^{*}$-algebras and their homomorphisms.
@@
Real $K$-theory $KO_{*}(A)$ is a $\mathbb{Z}$}-graded additive homology theory defined on the category $\mathscr{C}^{*}_{real}$ of all real $C^{*}$-algebras and taking values in $\mathbb{Z}$}-graded abelian groups. That is, it satisfies the following axioms:
@@
    and a long434343very long! It extends infinitely in both directions. exact sequence
@@
Complex $K$-theory $K_{*}(A)$ is a $\mathbb{Z}$}-graded additive homology theory defined on the category $\mathscr{C}^{*}_{cx}$ of all complex $C^{*}$-algebras and taking values in $\mathbb{Z}$}-graded abelian groups. That is, it satisfies the following axioms:
@@
We know from Cor. 5.9 that $KH_{*}$ is indeed a homology theory. It is periodic of period 4; the isomorphism is given by tensoring with the class of the identity in $KO_{0}({\mathbb{H}})$. In Remark 7.10 we noted that if $A$ is a real commutative $C^{*}$-algebra then there is a natural map
@@ 9. The Basics: Topological Spaces
A based topological space is a topological space $X$ together with a choice of basepoint $x_{o}\in X$. A based map $f:X\to Y$ of based topological spaces is a continuous map from $X$ to $Y$ with the property that $f(x_{0})=y_{0}$. We denote the set of based maps from $X$ to $Y$ by $F(X,Y)$ and based homotopy classes of based maps by $[X,Y]$. In particular, we may define the k’th homotopy group of $X$ by
@@
Note that on occasion we will want to deal with spaces and maps without specified basepoints, in which case we will denote the set of maps by $F(X,Y)^{u}$, where $u$ refers to “unbased.”  [^44] We considered using the term “baseless” but that seemed too derogatory. Following Bourbaki, we understand compact spaces to be compact and Hausdorff.
@@
Every locally compact space is compactly generated. So is every metrizable space, so this includes smooth manifolds and in particular all of the topological groups that we will need as well as the homogeneous spaces such as the space $O_{2k}/(O_{k}\times O_{k})$. All CW complexes, including spaces like $\mathbf{O}$ are compactly generated. If you confront a space $X$ that is not compactly generated then it is easy to retopologize it canonically to get a new compactly generated space $kX$ and a map $kX\to X$ which is the identity function and is continuous. [^45] See May [59] p. 37) for details.
@@
    then the pushout $Q$ is defined by taking the disjoint union of $X$ and $Y$ and identifying $f(w)$ with $g(w)$ for each $w\in W$.
@@ 10. The Basics: Fibrations, Fibre Bundles, Principal Bundles, and $\mathbf{BO}$
For each $b\in B$, the space $F_{b}=p^{-1}(b)$ is called the fibre over the point $b$.
@@
(cf. [59] p. 52) If the base space $B$ of a fibration $p:E\to B$ is path-connected then any two fibres $F_{x}$ and $F_{y}$ are homotopy equivalent.
@@
If $F$ and $B$ are CW-complexes then so is $\Omega B)$, and this implies by Whitehead’s theorem that they are homotopy equivalent. ∎
@@
A map $p:E\to B$ is a fibre bundle [^48] Some people spell the word “fiber” but we follow Norman Steenrod, the second author’s mathematical ancestor who, quite literally, wrote the (first!) book on “Fibre Bundles” [80] in 1951.
if it possesses a numerable open cover $\mathcal{U}$ and there is a fixed topological space $F$ and homeomorphisms $\phi_{U}$ for each open set $U$ in the cover such that the diagram
@@
    $G$ acts freely on $F_{b}$ ; that is, $xg=x$ implies that $g$ is the identity element of $G$.
@@
$B,BG]^{u}\longrightarrow Prin_{G}(B)$
from homotopy classes of maps $B\to BG$ to isomorphism classes of principal $G$-bundles over $B$. This map is an isomorphism (cf. Husemoller [44] 12.2 ). So this justifies calling $BG$ a ”classifying space” since homotopy classes of maps into it classify principal $G$-bundles.
@@
Define the associated bundle
@@
The associated bundle is a fibre bundle over $B$ with fibre $F$ and group $G$. Every fibre bundle with group $G$ arises in this manner.
@@
This implies by Proposition 10.7 that
@@
It turns out that $\mathbf{BO}$ has natural addition and multiplication maps which turn any $X,\mathbf{BO}]^{u}$ into a commutative semiring. Here is a sketch of the construction.
@@ 11. Bott Periodicity Revisited; First View of the Homotopy Road Map
$\Omega^{\mathbf{U}}\simeq\mathbf{U}.$
We need to take a much more careful look at the individual spaces $\Omega^{k}\mathbf{O}$ and $\Omega^{k}\mathbf{U}$. We are going to simply state the result that we need, and we refer the reader to Theorem 18.3 and to Part IV of Milnor, Morse Theory [62] for details.
@@
    The second column indicates description of the limiting Big Ten [^51] The second author grew up following Big Ten football and couldn’t resist the opportunity to call the ten spaces by their true name. homogeneous space in terms of the iterated loop space of the orthogonal group (or the unitary group, for the first two rows). Note that in the three cases $\mathbf{BU}\times{\mathbb{Z}}$, $\mathbf{BSp}\times{\mathbb{Z}}$, and $\mathbf{BO}\times{\mathbb{Z}}$, the space mentioned is the limit of the unstable space times the integers. So what is true, for example in the bottom row, is that
@@
    The third column gives the identification of the Big Ten speces in terms of iterated loop spaces of $\mathbf{U}$ and of $\mathbf{O}$ and illustrates their essential tie with Bott Periodicity. This will be explained in §17. For the moment we ask that it be taken on faith.
@@
    The fifth column indicates $\pi_{0}({\mathcal{R}}_{j})$. Remember that saying that $\pi_{0}=0$ means that the space is path connected. Saying that $\pi_{0}={\mathbb{Z}}_{2}$ says that the space has two connected components. If $\pi_{0}={\mathbb{Z}}$ then there are an infinite number of path components.
@@ 12. Next Contact with Physics: Low Homotopy Groups of the Ten-Fold Way
$0=\pi_{2}(G)\to\pi_{2}(G/H)\to\pi_{1}(H)\to\pi_{1}(G)\to\pi_{1}(G/H)\to\qquad(*)$
$\to\pi_{0}(H)\to\pi_{0}(G)\to\pi_{0}(G/H)\to 0$
The group $\pi_{2}(G)=0$ for all compact Lie groups as previously mentioned. If the group $H$ is path-connected [^55] A space $X$ is path-connected if $\pi_{0}(X)=0$. so that $\pi_{0}(H)=0$, then the sequence breaks apart into two simpler sequences-
$0\to\pi_{2}(G/H)\to\pi_{1}(H)\to\pi_{1}(G)\to\pi_{1}(G/H)\to 0\qquad(**)$
and
$0\to\pi_{0}(G)\to\pi_{0}(G/H)\to 0$
so that $\pi_{0}(G)\cong\pi_{0}(G/H)$.
@@ 1. The Groups $U_{n}$ and $SU_{n}$
These results connect up with the stable group ${\mathcal{R}}_{1}^{c}=\mathbf{U}$ since $\lim U_{n}=\mathbf{U}.$
@@ 2. The Homogenous Spaces $U_{n}/(U_{k}\times U_{n-k})$ and $SU_{n}/(SU_{k}\times SU_{n-k})$
$\phi=diag(A,B):U_{k}\times U_{n-k}\to U_{n}$
takes a pair $(A,B)$ of matrices and places them in the upper left and lower right parts of an $n\times n$ matrix (with zeros in the remaining two quarters). Sequence (**) applies, and so we have
@@
be given by $\delta(x)=(x,1)$. Then the map $\phi_{*}\delta_{*}(f)$ generates $\pi_{1}(U_{n})\cong{\mathbb{Z}}$ and hence $\phi_{*}$ is onto. This implies that
@@ 3. The Groups $O_{n}$ and $SO_{n}$
The groups $O_{n}$ have two path components, rotations and reflections, and $SO(n)$ is the closed, index $2$ subgroup of rotations. Thus
@@
These results tie to the group ${\mathcal{R}}_{1}=\mathbf{O}$, of course, since $\lim O_{n}\cong\mathbf{O}$.
@@ 4. The Homogeneous Spaces $O_{2n}/U_{n}$, $SO_{2n}/SU_{n}$, and $SO_{2n}/U_{n}$
denote the standard inclusion. The groups $U_{n}$ are path-connected for all $n$. So we have
$0\to\pi_{2}(O_{2n}/U_{n})\to\pi_{1}(U_{n})\overset{(\phi_{n})_{*}}{\to}\pi_{1}(O_{2n})\to\pi_{1}(O_{2n}/U_{n})\to 0\to\pi_{0}(O_{2n})\to\pi_{0}(O_{2n}/U_{n})\to 0.\qquad(**)$
@@
$\pi_{1}(O_{2n}/U_{n})=0\qquad and\qquad\pi_{2}(O_{2n}/U_{n})={\mathbb{Z}}$
@@
For $SO_{2n}/SU_{n}$ we apply sequence (**) since $SO_{2n}$ is path-connected. We get
@@
Finally, we consider $SO_{2n}/U_{n}$. The case $n=1$ is easy, since $U_{1}=SO_{2}$ and so all of the homotopy groups vanish. So assume that $n>1$. As $SO_{2n}$ is path-connected we may use sequence **. It reads
@@
To determine the map $\phi$, we can consider the commutative diagram
[…]
which implies that the map $\phi$ is onto. Thus
$\pi_{1}(SO_{2n}/U_{n})=0\qquad and\qquad\pi_{2}(SO_{2n}/U_{n})\cong{\mathbb{Z}}.$
@@
These results connect up with ${\mathcal{R}}_{2}=\lim O_{2n}/U_{n}$.
5. The Homogeneous Spaces $U_{n}/Sp_{n}$ and $SU_{n}/Sp_{n}$.
These are path-connected spaces. We have $\pi_{j}(Sp_{n})=0$ for $j\leq 2$ (see §7 ), so sequence (**) simplifies to
$0\to\pi_{2}(U_{n}/Sp_{n})\to 0\to{\mathbb{Z}}\to\pi_{1}(U_{n}/Sp_{n})\to 0.\qquad(**)$

### v3

@@ 8. Bott Periodicity
We now recap the principal properties of $KO_{*}$ for future use. Recall that $\mathscr{C}^{*}_{real}$ denotes the category of real $C^{*}$-algebras and their homomorphisms.
@@
Real $K$-theory $KO_{*}(A)$ is a $\mathbb{Z}$}-graded additive homology theory defined on the category $\mathscr{C}^{*}_{real}$ and taking values in $\mathbb{Z}$}-graded abelian groups. That is, it satisfies the following axioms:
@@
    and a long444444very long! It extends infinitely in both directions. exact sequence
@@
Complex $K$-theory $K_{*}(A)$ is a $\mathbb{Z}$}-graded additive homology theory defined on the category $\mathscr{C}^{*}_{cx}$ and taking values in $\mathbb{Z}$}-graded abelian groups. That is, it satisfies the following axioms:
@@
We know from Cor. 5.9 and Cor. 7.11 that $KH_{*}$ is indeed a homology theory. It is periodic of period 4; the isomorphism is given by tensoring with the class of the identity in $KO_{0}({\mathbb{H}})$. In Remark 7.10 we noted that if $A$ is a real commutative $C^{*}$-algebra then there is a natural map
@@ 9. The Basics: Topological Spaces
A based topological space $(X,x_{o})$ is a topological space $X$ together with a choice of basepoint $x_{o}\in X$. A based map
$f:(X,x_{o})\to(Y,y_{o})$
of based topological spaces (sometimes simply written $f:X\to Y$ when the context is clear) is a continuous map from $X$ to $Y$ with the property that $f(x_{0})=y_{0}$. We denote the set of based maps from $X$ to $Y$ by $F(X,Y)$ and based homotopy classes of based maps by $[X,Y]$. In particular, we may define the k’th homotopy group of $X$ by
@@
Note that on occasion we will want to deal with spaces and maps without specified basepoints, in which case we will denote the set of maps by $F(X,Y)^{ub}$, where $ub$ refers to “unbased.”  [^45] We considered using the term “baseless” but that seemed too derogatory. As noted previously, we follow Bourbaki and understand compact spaces to be compact and Hausdorff.
@@
Every locally compact space is compactly generated. So is every metrizable space, so this includes smooth manifolds and in particular all of the topological groups that we will need as well as the homogeneous spaces such as the space $O_{2k}/(O_{k}\times O_{k})$. All CW complexes, including spaces like $\mathbf{O}$ are compactly generated. If you confront a space $X$ that is not compactly generated then it is easy to retopologize it canonically to get a new compactly generated space $kX$ and a map $kX\to X$ which is the identity function and is continuous. [^46] See May ([59] p. 37) for details.
@@
    then the pushout $Q$ is defined by taking the disjoint union of $X$ and $Y$ modulo the equivalence relation generated by $f(w)\sim g(w)$ for each $w\in W$.
@@ 10. The Basics: Fibrations, Fibre Bundles, Principal Bundles, and $\mathbf{BO}$
For each $b\in B$, the space $F_{b}=p^{-1}(b)$ is called the fibre [^48] Some people spell the word “fiber” but we follow Norman Steenrod, the second author’s mathematical ancestor who, quite literally, wrote the (first!) book on “Fibre Bundles” [80] in 1951. over the point $b$.
@@
(cf. [59] p. 52) If the base space $B$ of a fibration $p:E\to B$ is path-connected [^50] A space $X$ is path-connected if given any two points $x_{0},\,x_{1}\in X$ there is a continuous path $x_{t}$ in $X$ connecting them. This is equivalent to asserting that $\pi_{0}(X)=0$. then any two fibres $F_{x}$ and $F_{y}$ are homotopy equivalent.
@@
If $F$ and $B$ are CW-complexes then so is $\Omega B$, and this implies by Whitehead’s theorem that they are homotopy equivalent. ∎
@@
A map $p:E\to B$ is a fibre bundle if it possesses a numerable open cover $\mathcal{U}$ and there is a fixed topological space $F$ and homeomorphisms $\phi_{U}$ for each open set $U$ in the cover such that the diagram
@@
    $G$ acts freely on $F_{b}$ ; that is, $xg=x$ for all $x\in X$ implies that $g$ is the identity element of $G$.
@@
$B,BG]^{ub}\longrightarrow Prin_{G}(B)$
from homotopy classes of maps $B\to BG$ to isomorphism classes of principal $G$-bundles over $B$. This map is an isomorphism (cf. Husemoller [44] 12.2 ). So this justifies calling $BG$ a “classifying space” since homotopy classes of maps into it classify principal $G$-bundles.
@@
Define the associated bundle by
@@
The associated bundle is a fibre bundle
$F\longrightarrow E\times_{G}F\longrightarrow B$
with group $G$. Every fibre bundle with group $G$ arises up to isomorphism in this manner.
@@
Proposition 10.7 tells us that
@@
It turns out that $\mathbf{BO}$ has natural addition and multiplication maps which turn any $X,\mathbf{BO}]^{ub}$ into a commutative semiring. Here is a sketch of the construction.
@@ 11. Bott Periodicity Revisited; First View of the Homotopy Road Map
$\Omega^{2}\mathbf{U}\simeq\mathbf{U}.$
We need to take a much more careful look at the individual spaces $\Omega^{k}\mathbf{O}$ and $\Omega^{k}\mathbf{U}$. We shall simply state the result that we need, and we refer the reader to Theorem 18.3 and to Part IV of Milnor, Morse Theory [62] for details.
@@
    The second column is the direct limit of the first column. It indicates description of the limiting Big Ten [^53] The second author grew up following Big Ten football and couldn’t resist the opportunity to call the ten spaces by their true name. homogeneous space in terms of the iterated loop space of the orthogonal group (or the unitary group, for the first two rows). Note that in the three cases $\mathbf{BU}\times{\mathbb{Z}}$, $\mathbf{BSp}\times{\mathbb{Z}}$, and $\mathbf{BO}\times{\mathbb{Z}}$, the space mentioned is the limit of the unstable space times the integers. So what is true, for example in the bottom row, is that
@@
    The third column gives the identification of the Big Ten spaces in terms of iterated loop spaces of $\mathbf{U}$ and of $\mathbf{O}$ and illustrates their essential tie with Bott Periodicity. This will be explained in §17. For the moment we ask that it be taken on faith.
@@
    The fifth column indicates $\pi_{0}({\mathcal{R}}_{j})$. Remember that saying that $\pi_{0}=0$ means that the space is path-connected. Saying that $\pi_{0}={\mathbb{Z}}_{2}$ says that the space has two connected components. If $\pi_{0}={\mathbb{Z}}$ then there are an infinite number of path components.
@@
Thus
$\pi_{5}(O_{k})=\pi_{0}(\Omega^{5}O_{k})=0\qquad k\geq 7$
and hence the eighth row lists $\pi_{0}(\Omega^{5}(\mathbf{O})=0$ even though
$\pi_{5}(O_{3})\neq 0\qquad\pi_{5}(O_{4})\neq 0\qquad\pi_{5}(O_{5})\neq 0\qquad\pi_{5}(O_{6})\neq 0.$
@@ 12. Next Contact with Physics: Low Homotopy Groups of the Ten-Fold Way
$0=\pi_{2}(G)\to\pi_{2}(G/H)\to\pi_{1}(H)\to\pi_{1}(G)\to\pi_{1}(G/H)\to\\ \to\pi_{0}(H)\to\pi_{0}(G)\to\pi_{0}(G/H)\to 0$ (12.1)
The group $\pi_{2}(G)=0$ for all compact Lie groups as previously mentioned. If the group $H$ is path-connected so that $\pi_{0}(H)=0$, then the sequence breaks apart into two simpler sequences-
$\begin{split}0\to\pi_{2}(G/H)&\to\pi_{1}(H)\to\pi_{1}(G)\to\pi_{1}(G/H)\to 0\;,\\ \text{and}&0\to\pi_{0}(G)\to\pi_{0}(G/H)\to 0\;.\end{split}$ (12.2)
In particular, there is an isomorphism $\pi_{0}(G)\cong\pi_{0}(G/H)$.
@@ 1. The Groups $U_{n}$ and $SU_{n}$
[…]
@@ 2. The Homogenous Spaces $U_{n}/(U_{k}\times U_{n-k})$ and $SU_{n}/(SU_{k}\times SU_{n-k})$
$\phi=\mathrm{diag}(A,B):U_{k}\times U_{n-k}\to U_{n}$
takes a pair $(A,B)$ of matrices and places them in the upper left and lower right parts of an $n\times n$ matrix (with zeros in the remaining two quarters). Sequence (12.2) applies, and so we have
@@
be given by $\delta(x)=(x,1)$. Then the map $\phi_{*}\delta_{*}(f)$ generates $\pi_{1}(U_{n})\cong{\mathbb{Z}}$ and hence $\phi_{*}$ is surjective. This implies that
@@ 3. The Groups $O_{n}$ and $SO_{n}$
The groups $O_{n}$ have two path components, rotations and reflections, and $SO_{n}$ is the closed, index $2$ subgroup of rotations. Thus
@@
84
[…]
@@ 4. The Homogeneous Spaces $O_{2n}/U_{n}$, $SO_{2n}/SU_{n}$, and $SO_{2n}/U_{n}$
denote the standard inclusion. The groups $U_{n}$ are path-connected for all $n$. So we have from Sequence (12.2),
$\begin{split}0\to\pi_{2}(O_{2n}/U_{n})&\to\pi_{1}(U_{n})\overset{(\phi_{n})_{*}}{\to}\pi_{1}(O_{2n})\to\pi_{1}(O_{2n}/U_{n})\to 0\\ \text{and~}&0\to\pi_{0}(O_{2n})\to\pi_{0}(O_{2n}/U_{n})\to 0.\end{split}$
@@
$\pi_{1}(O_{2n}/U_{n})=0\qquad\text{and}\qquad\pi_{2}(O_{2n}/U_{n})={\mathbb{Z}}$
@@
For $SO_{2n}/SU_{n}$ we apply Sequence (12.2) since $SO_{2n}$ is path-connected. We get
@@
Finally, we consider $SO_{2n}/U_{n}$. The case $n=1$ is easy, since $U_{1}=SO_{2}$ and so all of the homotopy groups vanish. So assume that $n>1$. As $SO_{2n}$ is path-connected we may use Sequence (12.2). It reads
@@
To determine the map $\phi_{*}$, we can consider the commutative diagram
${\pi_{1}(U_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}\scriptstyle{\cong}$${\pi_{1}(SO_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}\textstyle{{\mathbb{Z}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(U_{n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{*}}$${\pi_{1}(SO_{2n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${{\mathbb{Z}}_{2}}$
which implies that the map $\phi_{*}$ is surjective. Thus
$\pi_{1}(SO_{2n}/U_{n})=0\qquad\text{and}\qquad\pi_{2}(SO_{2n}/U_{n})\cong{\mathbb{Z}}.$
@@
[…]
5. The Homogeneous Spaces $U_{n}/Sp_{n}$ and $SU_{n}/Sp_{n}$
These are path-connected spaces. We have $\pi_{j}(Sp_{n})=0$ for $j\leq 2$ (see §7), so Sequence (12.2) simplifies to
$0\to\pi_{2}(U_{n}/Sp_{n})\to 0\to{\mathbb{Z}}\to\pi_{1}(U_{n}/Sp_{n})\to 0\;.$